OpenStudy (anonymous):

Express cos(2x) in terms of sin(x).

OpenStudy (anonymous):

This involves an trigonometric identity! Which one involves 2x?

OpenStudy (anonymous):

This is the double angle for cosine, but I don't know how to experss it in terms of sin(x).

OpenStudy (anonymous):

\(\ \large \text{Okay, let's go back... What's the double angle identity for cosine?} \)

OpenStudy (anonymous):

1-2sin^2(x) = cos(2x)

OpenStudy (anonymous):

\(\ \Large \text{Exactly! That's it! That's in terms of sin(x)!} \)

OpenStudy (anonymous):

Oh! Okay :) Thanks. Do you mind helping me with another trig problem?

OpenStudy (anonymous):

\(\ \Huge \text {I'll try my best ;)} \)

OpenStudy (anonymous):

\(\ \text {By the way -- Welcome to OpenStudy!} \)

OpenStudy (anonymous):

Express  and find the sum of a+b+c= ? without using a calculator

OpenStudy (anonymous):

that's radical of a+b(radical c)

OpenStudy (anonymous):

and thanks :)

OpenStudy (anonymous):

Hmmm...I'm not really sure. I'd say start by knowing what cos60 is and working from there?

OpenStudy (anonymous):

Oh, alright thanks anyway :)

OpenStudy (anonymous):

1-2 sin^2 x

OpenStudy (anonymous):

we know that cos 2a =cos^2a-sin^2a =1-sin^2a-sin^2a=1-2sin^2a

OpenStudy (anonymous):

i think u got this